# GNSS-based Lunar Orbit and Clock Estimation With Stochastic Cloning UD Filter

### v1

@@ 1 Introduction
Lunar exploration has been a topic of significant interest in recent years, driven by international efforts to establish a sustainable presence on the Moon. To meet the navigation and timing needs of these missions, NASA, ESA, and JAXA have proposed the Lunar Augmented Navigation Service (LANS) as part of LunaNet [Israel2020], the lunar counterpart to Earth’s Global Navigation Satellite System (GNSS). The LANS aims to provide precise positioning, navigation, and timing (PNT) services to assets on the lunar surface and in orbit by leveraging a constellation of lunar satellites that broadcast Augmented Forward Signals (AFS) [LNISv5]. Specifically, the National Aeronautics and Space Administration (NASA), the European Space Agency (ESA), and the Japan Aerospace Exploration Agency (JAXA) plan to deploy the Lunar Communications Relay and Navigation Systems (LCRNS) [gramling2024], the Lunar Communications and Navigation Service (LCNS) [traveset2024], and the Lunar Navigation Satellite System (LNSS) [Murata2022], respectively.
@@
Initial research on terrestrial GNSS-based lunar orbit and clock estimation proposed both snapshot least-squares and Extended Kalman Filtering (EKF) approaches to estimate satellite states using pseudorange measurements [Capuano2016, Capuano2017], and were verified for feasibility through simulations. Recently, the Lunar GNSS Receiver Experiment (LuGRE) mission demonstrated this concept in flight [parker2022lugre, Konitzer2022LuGRE]. The mission successfully tracked terrestrial GNSS signals from lunar orbit and performed onboard estimation using pseudorange measurements. However, the resulting positioning accuracy was limited to the order of kilometers [insidegnss2025lugre]. This limitation stemmed primarily from the short duration of the experiment, which prevented the accumulation of sufficient data. Due to the poor geometry of GNSS satellites at lunar distances, the filter requires longer durations to converge to a stable solution. The reliance on pseudorange measurements further restricts accuracy due to their relatively high measurement noise compared to carrier phase measurements.
@@
does not hold for measurements $Z_{k}$ that depend on both the current state $X_{k}$ and a past state $X_{k-1}$. To handle this, [Iiyama2024TDCPNavigation] employed stochastic cloning [Roumeliotis2002StochasticCloning], while [Mina2025LCRNS] utilized a delayed-state EKF. These studies demonstrated that incorporating TDCP measurements improves orbit and clock estimation accuracy by 10–30% compared to pseudorange-only solutions, achieving sub-10 meter accuracy in simulation.
Despite these advancements, existing methods suffer from several limitations. First, the Kalman filters used in prior works risk divergence due to the low observability of the system, particularly when processing very precise TDCP measurements. In such scenarios, the covariance matrix in standard Kalman filters can lose its positive semi-definiteness (PSD) due to numerical errors, leading to filter divergence and degraded performance [Tapley2004]. In aerospace applications, it is standard practice to use square-root formulations [Potter1963StatFiltering] or UD-factorization-based Kalman filters [Thornton1976TriangularCovariance, BiermanThornton1977_KalmanComparison] to mitigate this. Square-root filters were notably used in the Apollo program to maximize numerical precision within 8-bit limits. Alternatively, the UD-filter avoids square-root operations, making it computationally more efficient, and has been recently adopted in the Orion Navigation System [Sud2014OrionEFT1Nav]. These methods replace the covariance matrix with its factorization, ensuring symmetry and positive semi-definiteness by construction [DSouzaZanetti2019]. However, to the best of the authors’ knowledge, variants of these covariance-factorization filters capable of processing delayed-state measurements such as TDCP have not yet been presented in the literature.
@@
Third, the effects of general and special relativity on GNSS signals and the local clocks of lunar satellites have not been adequately addressed in prior ODTS works. Relativistic effects induce coupling between orbital states and clock bias states, leading to systematic biases in measurements and estimated clock biases if not properly modeled [seyffert2025]. While the LuGRE mission synchronized its clock to GPS Time (GPST) [Konitzer2022LuGRE], future lunar satellites aiming for high-precision PNT services will need to synchronize clocks to lunar-based reference time frames [Bourgoin_2025], making accurate modeling of these relativistic effects essential.
@@ 2.1 State Definition
$X_{k}=\begin{bmatrix}\mathbf{r}_{k}^{\mathrm{LCRS}}&\mathbf{v}_{k}^{\mathrm{LCRS}}&c\,\delta t_{k}&c\,\delta\dot{t}_{k}&c\,\delta\ddot{t}_{k}&\gamma_{\mathrm{srp}}\end{bmatrix}^{\!T},$ (2)
where $\mathbf{r}_{k}^{\mathrm{LCRS}}$ and $\mathbf{v}_{k}^{\mathrm{LCRS}}$ denote the position and velocity of the lunar satellite expressed in the Lunar Celestial Reference System (LCRS). The LCRS is centered at the Moon’s center of mass and shares its orientation with the Barycentric Celestial Reference System (BCRS) and Geocentric Celestial Reference System (GCRS). For notational brevity, the superscript $\mathrm{LCRS}$ is omitted in subsequent sections unless other reference frames are explicitly indicated.
The remaining components of the state vector include the clock bias $\delta t_{k}$, clock drift $\delta\dot{t}_{k}$, and clock acceleration $\delta\ddot{t}_{k}$, all scaled by the speed of light $c$, as well as the solar radiation pressure (SRP) coefficient $\gamma_{\mathrm{srp}}$. The clock states are defined with respect to Lunar Coordinate Time (TCL), which serves as the coordinate time of the LCRS. A detailed discussion of TCL and its relationship to other time scales is provided in Section 2.3.
@@ 2.2 Orbit Dynamics
$\begin{bmatrix}\dot{\mathbf{r}}_{k}\\ \dot{\mathbf{v}}_{k}\end{bmatrix}$ $=\begin{bmatrix}\mathbf{v}_{k}\\ \mathbf{a}_{g}(t_{k},\mathbf{r}_{k})+\mathbf{a}_{\mathrm{srp}}(t_{k},\mathbf{r}_{k},\gamma_{\mathrm{srp}})\end{bmatrix}+\boldsymbol{\epsilon}^{rv}_{k},$ (3)
where $\boldsymbol{\epsilon}^{rv}_{k}$ represents the process noise associated with unmodeled dynamics and modeling uncertainties.
@@
$\mathbf{r}^{\mathrm{TCL}}$ $=\mathbf{r}^{\mathrm{TCB}}=\frac{1}{1-L_{B}}\,\mathbf{r}^{\mathrm{TDB}},$ (6)
@@
where $L_{B}=1.550519768\times 10^{-8}$ is the defining constant relating TCB and TDB. This conversion is essential when utilizing planetary ephemerides and gravitational parameters provided in TDB-compatible form.
@@ 2.3.1 Lunar Coordinate Time (TCL) and Lunar Time (LT)
LANS satellites are expected to broadcast a common lunar time scale for navigation and timing. Although the final definition has not yet been standardized, we assume this reference to be Lunar Time (LT), defined to remain close to the proper time of clocks located on the selenoid. Before introducing LT, we first review the definition of Lunar Coordinate Time (TCL), following [Kopeikin_2024].
TCL is the relativistic coordinate time associated with the Lunar Celestial Reference System (LCRS), analogous to Geocentric Coordinate Time (TCG) for the Geocentric Celestial Reference System (GCRS) and Barycentric Coordinate Time (TCB) for the Barycentric Celestial Reference System (BCRS). The transformation between TCL and TCB is defined by accounting for the Moon’s barycentric motion and the external gravitational potential acting on the Moon:
$\frac{d\mathrm{TCL}}{d\mathrm{TCB}}=1-\frac{1}{c^{2}}\left(\frac{v_{m}^{2}}{2}+U_{\mathrm{ext}}(\mathbf{x}_{m})\right)+\mathcal{O}\!\left(\frac{1}{c^{4}}\right),$ (8)
where $\mathbf{x}_{m}$ and $v_{m}$ denote the Moon’s barycentric position and velocity, respectively, and $U_{\mathrm{ext}}(\mathbf{x}_{m})$ represents the gravitational potential of all bodies excluding the Moon, evaluated at the Moon’s location.
@@
Using these relations, terrestrial GNSS time (GPST) can be transformed into LT via the sequence [IERS2010, Kopeikin_2024]:
@@
where $L_{G}=6.969290134\times 10^{-10}$ and $L_{L}=3.14027\times 10^{-11}$ are linear scale constants relating TCG to TT and TCL to LT, respectively. The term $\Delta(\mathrm{TCG}-\mathrm{TCL})$ is given by (9). The coefficient $L_{L}=\Phi_{L}/c^{2}$ is derived from the reference lunar potential $\Phi_{L}=2.822336927\times 10^{6}~\mathrm{m}^{2}/\mathrm{s}^{2}$ [Kopeikin_2024], though its final value remains under international standardization.
Efficient onboard implementation of the TCL–TCG transformation requires compact analytical approximations of $\Delta(\mathrm{TCG}-\mathrm{TCL})$. While global polynomial approximations are challenging to maintain over long durations [Lu2025_LTE440], piecewise polynomial or spline-based representations over short intervals may offer a practical compromise between accuracy and computational efficiency.
@@ 2.3.3 Shapiro Time Delay
where $\mu_{S}$ is the solar gravitational parameter, $r_{k}^{\mathrm{BCRF}}$ and $r_{tx,i}^{\mathrm{BCRF}}$ are the barycentric distances of the receiver and transmitter from the Sun, respectively, and $R_{k,i}$ is the Euclidean distance between them.
@@ 2.4.1 Pseudorange and Carrier-Phase Measurements
where $\bar{\rho}_{i}(t_{k})$ is the nominal geometric-plus-clock range (defined below), $I_{i,L}^{P}$ and $I_{i,L}^{C}$ denote ionospheric/plasmaspheric contributions to code and carrier phase, respectively, $\Delta t^{S}_{k,i}$ is the Shapiro delay expressed in time, $\delta P_{i,L}$ and $\delta C_{i,L}$ collect instrumental and modeling biases, and $w_{\rho,k}$ and $w_{\Phi,k}$ represent measurement noise.
@@
$\bar{\rho}_{i}(t_{k})$ $=\left\|\mathbf{r}_{k}^{\mathrm{ECEF}}-\bar{\mathbf{r}}_{tx,i}^{\mathrm{ECEF}}(\bar{t}_{tx,i},\boldsymbol{\alpha})\right\|+c\left[\delta t_{k}^{\mathrm{GPST}}-\bar{\delta t}_{tx,i}^{\mathrm{GPST}}(\bar{t}_{tx,i},\boldsymbol{\alpha})\right],$ (17)
@@
where $\mathbf{r}_{k}^{\mathrm{ECEF}}$ is the receiver position in ECEF at $t_{k}$, $\bar{\mathbf{r}}_{tx,i}^{\mathrm{ECEF}}(\cdot)$ and $\bar{\delta t}_{tx,i}^{\mathrm{GPST}}(\cdot)$ are the broadcast-ephemeris transmitter position and clock bias evaluated at the estimated transmission time $\bar{t}_{tx,i}$ (Sec. 2.4.3), and $\boldsymbol{\alpha}$ denotes the broadcast ephemeris parameters. All symbols are summarized in Table 1.
@@
$\mathbf{r}_{k}^{ECEF}$ Receiver position in Earth-Centered Earth-Fixed (ECEF) at time step $k$
$\bar{\mathbf{r}}_{tx,i}^{ECEF}(\bar{t}_{tx,i},\boldsymbol{\alpha})$ Estimated transmitter satellite position in ECEF at transmission time
@@
$\bar{\delta}t_{tx,i}^{GPST}(\bar{t}_{tx,i},\boldsymbol{\alpha})$ Estimated transmitter satellite clock bias in GPS Time (GPST) at transmission time

### v2

@@ 1 Introduction
Lunar exploration has been a topic of significant interest in recent years, driven by international efforts to establish a sustainable presence on the Moon. To meet the navigation and timing needs of these missions, the National Aeronautics and Space Administration (NASA), European Space Agency (ESA), and Japan Aerospace Exploration Agency (JAXA) have proposed the Lunar Augmented Navigation Service (LANS) as part of LunaNet [Israel2020], the lunar counterpart to Earth’s Global Navigation Satellite System (GNSS). The LANS aims to provide precise positioning, navigation, and timing (PNT) services to assets on the lunar surface and in orbit by leveraging a constellation of lunar satellites that broadcast Augmented Forward Signals (AFS) [LNISv5]. Specifically, the National Aeronautics and Space Administration (NASA), the European Space Agency (ESA), and the Japan Aerospace Exploration Agency (JAXA) plan to deploy the Lunar Communications Relay and Navigation Systems (LCRNS) [gramling2024], the Lunar Communications and Navigation Service (LCNS) [traveset2024], and the Lunar Navigation Satellite System (LNSS) [Murata2022], respectively.
@@
Initial research on terrestrial GNSS-based lunar orbit and clock estimation proposed both snapshot least-squares and Extended Kalman Filtering (EKF) approaches to estimate satellite states using pseudorange measurements [Capuano2014EarthMoonGNSS, Capuano2017], and were verified for feasibility through simulations. Recently, the Lunar GNSS Receiver Experiment (LuGRE) mission demonstrated this concept in flight [parker2022lugre, Konitzer2022LuGRE]. The mission successfully tracked terrestrial GNSS signals from lunar orbit and performed onboard estimation using pseudorange measurements. However, the resulting positioning accuracy was limited to the order of kilometers [insidegnss2025lugre]. This limitation stemmed primarily from the short duration of the experiment (one-hour observation window), and the relatively low stability of the clock compared to navigation satellites (Voltage Controlled Temperature Compensated Crystal Oscillator (VCTCXO) [Konitzer2022LuGRE]), which prevented the accumulation of sufficient data. Due to the poor geometry of GNSS satellites at lunar distances, the filter requires longer durations to converge to a stable solution. The reliance on pseudorange measurements further limits accuracy because their noise is higher than that of carrier-phase measurements.
@@
does not hold for measurements $Z_{k}$ that depend on both the current state $X_{k}$ and a past state $X_{k-1}$. To handle this, [Iiyama2024TDCPNavigation] employed stochastic cloning, a method that was originally proposed in [Roumeliotis2002StochasticCloning], while [Mina2025LCRNS] utilized a delayed-state EKF. These studies demonstrated that incorporating TDCP measurements improves orbit and clock estimation accuracy by 10–30% compared to pseudorange-only solutions, achieving sub-10 meter accuracy in simulation.
Despite these advancements, existing methods remain limited. First, the Kalman filters used in prior works risk divergence due to the low observability of the system, particularly when processing very precise TDCP measurements. In such scenarios, the covariance matrix in standard Kalman filters can lose its positive semi-definiteness (PSD) due to numerical errors, leading to filter divergence and degraded performance [Tapley2004]. In aerospace applications, it is standard practice to use square-root formulations [Potter1963StatFiltering] or UD-factorization-based Kalman filters [Thornton1976TriangularCovariance, BiermanThornton1977_KalmanComparison] to mitigate this. Square-root filters were notably used in the Apollo program to maximize numerical precision within 8-bit limits [Fraser2019ApolloMemories]. Alternatively, the UD-filter factorizes the covariance matrix as $P=UDU^{\top}$, where $U$ is a upper triangular matrix with ones on diagonals and $D$ is a diagonal matrix. UD-filter avoids square-root operations, making it computationally more efficient, and has been recently adopted in the Orion Navigation System [Sud2014OrionEFT1Nav]. These methods replace the covariance matrix with its factorization, ensuring symmetry and positive semi-definiteness by construction [DSouzaZanetti2019]. However, to the best of the authors’ knowledge, variants of these covariance-factorization filters capable of processing delayed-state measurements such as TDCP have not yet been presented in the literature.
@@
Third, the effects of general and special relativity on GNSS signals and the local clocks of lunar satellites have not been adequately addressed in prior GNSS-based lunar orbit determination and clock synchronization (ODTS) works. Relativistic effects induce coupling between orbital states and clock bias states, leading to systematic biases in measurements and estimated clock biases if not properly modeled [seyffert2025]. While the LuGRE mission synchronized its clock to GPS Time (GPST) [Konitzer2022LuGRE], future lunar satellites aiming for high-precision PNT services will need to synchronize clocks to lunar-based reference time frames [Bourgoin_2025, Turyshev2026RelativisticTimescales, Kopeikin_2024], making accurate modeling of these relativistic effects essential.
@@ 2.1 State Definition
$X_{k}=\begin{bmatrix}\mathbf{r}_{k}^{\mathrm{LCRF}}&\mathbf{v}_{k}^{\mathrm{LCRF}}&c\,\delta t_{k}&c\,\delta\dot{t}_{k}&c\,\delta\ddot{t}_{k}&\gamma_{\mathrm{srp}}\end{bmatrix}^{\!T},$ (2)
where $\mathbf{r}_{k}^{\mathrm{LCRF}}$ and $\mathbf{v}_{k}^{\mathrm{LCRF}}$ denote the position and velocity of the lunar satellite expressed in the Lunar Celestial Reference Frame (LCRF), which is the realization of the Lunar Celestial Reference System (LCRS). The LCRF is centered at the Moon’s center of mass and shares its orientation with the Barycentric Celestial Reference Frame (BCRF) and Geocentric Celestial Reference Frame (GCRF). For notational brevity, the superscript $\mathrm{LCRF}$ is omitted in subsequent sections unless other reference frames are explicitly indicated.
The remaining components of the state vector include the clock bias $\delta t_{k}$, clock drift $\delta\dot{t}_{k}$, and clock acceleration $\delta\ddot{t}_{k}$, all scaled by the speed of light $c$, as well as the solar radiation pressure (SRP) coefficient $\gamma_{\mathrm{srp}}$. The clock states are defined with respect to Lunar Coordinate Time (TCL), which serves as the coordinate time of the LCRF. A detailed discussion of TCL and its relationship to other time scales is provided in Section 2.3.
@@ 2.2 Orbit Dynamics
$\begin{bmatrix}\dot{\mathbf{r}}_{k}\\ \dot{\mathbf{v}}_{k}\end{bmatrix}$ $=\begin{bmatrix}\mathbf{v}_{k}\\ \mathbf{a}_{g}(t_{k},\mathbf{r}_{k})+\mathbf{a}_{\mathrm{srp}}(t_{k},\mathbf{r}_{k},\gamma_{\mathrm{srp}})\end{bmatrix}$ (3)
@@
$\mathcal{X}^{\mathrm{TCL}}$ $=\mathcal{X}^{\mathrm{TCB}}=\frac{1}{1-L_{B}}\,\mathbf{X}^{\mathrm{TDB}},$ (6)
@@
where $\mathcal{X}$ denotes the scale for positions, and $L_{B}=1.550519768\times 10^{-8}$ is the defining constant relating TCB and TDB. This conversion is essential when utilizing planetary ephemerides and gravitational parameters provided in TDB-compatible form.
@@ 2.3.1 Lunar Coordinate Time (TCL) and Lunar Time (LT)
LANS satellites are expected to broadcast a common lunar time scale for navigation and timing. Although the final definition has not yet been standardized, we assume this reference to be Lunar Time (LT), defined to remain close to the proper time of clocks located on the selenoid, as assumed in [Bourgoin_2025, Turyshev2026RelativisticTimescales]. Before introducing LT, we first review the definition of TCL, following [Kopeikin_2024].
TCL is the relativistic coordinate time associated with the LCRS, analogous to Geocentric Coordinate Time (TCG) for the GCRS and Barycentric Coordinate Time (TCB) for the BCRS. The transformation between TCL and TCB is defined by accounting for the Moon’s barycentric motion and the external gravitational potential acting on the Moon:
$\frac{d\mathrm{TCL}}{d\mathrm{TCB}}=1-\frac{1}{c^{2}}\left(\frac{\|\mathbf{v}_{m}\|^{2}}{2}+U_{\mathrm{ext}}(\mathbf{x}_{m})\right)+\mathcal{O}\!\left(\frac{1}{c^{4}}\right),$ (8)
where $\mathbf{x}_{m}$ and $\mathbf{v}_{m}$ denote the Moon’s barycentric position and velocity, respectively, and $U_{\mathrm{ext}}(\mathbf{x}_{m})$ represents the gravitational potential of all bodies excluding the Moon, evaluated at the Moon’s location.
@@
Using these relations, terrestrial GNSS time (GPST) can be transformed into LT via the sequence [IERS2010, Kopeikin_2024, Turyshev2026RelativisticTimescales]:
@@
where $L_{G}=6.969290134\times 10^{-10}$ and $L_{L}$ are linear scale constants relating TCG to TT and TCL to LT, respectively, and $T_{0}=2443144.5003725$ JD (Julian Date) is the reference epoch. The term $\Delta(\mathrm{TCG}-\mathrm{TCL})$ is given by (9). The coefficient $L_{L}=\Phi_{L}/c^{2}$ is depends on the reference chosen lunar potential $\Phi_{L}$. Reference [Kopeikin_2024] adopts a lunar gravitational potential of $\Phi_{L}=2.822336927\times 10^{6},\mathrm{m}^{2}/\mathrm{s}^{2}$, yielding $L_{L}=3.14027\times 10^{-11}$, whereas [Turyshev2026RelativisticTimescales] assumes $\Phi_{L}=2.82123744381\times 10^{6},\mathrm{m}^{2}/\mathrm{s}^{2}$, resulting in $L_{L}=3.13905\times 10^{-11}$. The definitive value of $L_{L}$ has not yet been established through international standardization, and the initial lunar timescale offset, $T_{L0}$, likewise remains undefined.
Efficient onboard implementation of the TCL–TCG transformation requires compact analytical approximations of $\Delta(\mathrm{TCG}-\mathrm{TCL})$. While global polynomial approximations are challenging to maintain over long durations [Lu2025_LTE440], piecewise polynomial or spline-based representations over short intervals may offer a practical compromise between accuracy and computational efficiency. It would also be necessary to represent TCL–TCG in TCL (Eq (9) is in TCG), to execute inverse transformation without root-solving.
@@ 2.3.3 Shapiro Time Delay
where $\mu_{S}$ is the solar gravitational parameter, $r_{k}^{\mathrm{BCRF}}$ and $r_{tx,i}^{\mathrm{BCRF}}$ are the barycentric distances of the receiver and transmitter from the Sun, respectively, and $R_{k,i}$ is the Euclidean distance between them. For Earth-Moon links, the Shapiro delays from the Sun is in $20\text{\,}\mathrm{n}\mathrm{s}$ to $30\text{\,}\mathrm{n}\mathrm{s}$, which dominates the delay when compared with $0.1\text{\,}\mathrm{n}\mathrm{s}$ to $0.2\text{\,}\mathrm{n}\mathrm{s}$ from Earth, and $1\text{\,}\mathrm{p}\mathrm{s}$ to $3\text{\,}\mathrm{p}\mathrm{s}$ from the Moon [Turyshev2026RelativisticTimescales].
@@ 2.4.1 Pseudorange and Carrier-Phase Measurements
where $\bar{\rho}_{i}(t_{k})$ is the nominal geometric-plus-clock range (defined below), $I_{i,L}^{P}$ and $I_{i,L}^{C}$ denote ionospheric/plasmaspheric contributions to code and carrier phase, respectively, $\Delta t^{S}_{k,i}$ is the Shapiro delay expressed in time, $N_{i,L}$ is the integer ambiguity, $\delta P_{i,L}$ and $\delta C_{i,L}$ collect instrumental and modeling biases, and $w_{\rho,k}$ and $w_{\Phi,k}$ represent measurement noise.
@@
$\bar{\rho}_{i}(t_{k})$ $=\left\|\mathbf{r}_{k}^{\mathrm{ITRF}}-\bar{\mathbf{r}}_{tx,i}^{\mathrm{ITRF}}(\bar{t}_{tx,i},\boldsymbol{\alpha})\right\|+c\left[\delta t_{k}^{\mathrm{GPST}}-\bar{\delta t}_{tx,i}^{\mathrm{GPST}}(\bar{t}_{tx,i},\boldsymbol{\alpha})\right],$ (17)
@@
where $\mathbf{r}_{k}^{\mathrm{ITRF}}$ is the receiver position in International Terrestrial Reference Frame (ITRF) at $t_{k}$, $\bar{\mathbf{r}}_{tx,i}^{\mathrm{ITRF}}(\cdot)$ and $\bar{\delta t}_{tx,i}^{\mathrm{GPST}}(\cdot)$ are the broadcast-ephemeris transmitter position and clock bias evaluated at the estimated transmission time $\bar{t}_{tx,i}$ (Sec. 2.4.3), and $\boldsymbol{\alpha}$ denotes the broadcast ephemeris parameters. All symbols are summarized in Table 1.
@@
$\mathbf{r}_{k}^{ITRF}$ Receiver position in International Terrestrial Reference Frame (ITRF) at time step $k$
$\bar{\mathbf{r}}_{tx,i}^{ITRF}(\bar{t}_{tx,i},\boldsymbol{\alpha})$ Estimated transmitter satellite position in ITRF at transmission time
@@
$\bar{\delta}t_{tx,i}^{GPST}(\bar{t}_{tx,i},\boldsymbol{\alpha})$ Estimated transmitter satellite clock bias in GPST at transmission time